Four years ago Tarick was three years older than twice Deonte's age then. Six years from now the sum of their ages will be 39. Find the difference in their present ages.
*****I have tried solving this word problem several times but I am not getting the correct answer. (I already know the answer, I just need a little help solving it) My teacher gave the answer 25/3. I can't solve it to get that answer. I think I'm just not setting it up correctly or I'm confused by the wording of the question.*****

Answer by scott8148(6628)

"Four years ago Tarick was three years older than twice Deonte's age then" ___ T-4 = 2(D-4) + 3 ___ T-4 = 2D-5

"Six years from now the sum of their ages will be 39" ___ T+6 + D+6 = 39 ___ T + D = 27 ___ T = 27-D

substituting ___ (27-D)-4 = 2D-5 ___ 28 = 3D ___ 28/3 = D

substituting ___ T = 27-(28/3) ___ T = 53/3

T-D = (53/3)-(28/3) = 25/3